OpenStudy (anonymous):

9x^2-6x+1=0 a) solve b) find x-intercepts

OpenStudy (anonymous):

3x-1^2

OpenStudy (anonymous):

(3x-1)(3x-1)

OpenStudy (anonymous):

x=-3/1

OpenStudy (anonymous):

what are the solutions? what are the intercepts?

OpenStudy (anonymous):

9(-3)-6(-3)+1=y intercept

OpenStudy (anonymous):

if everything was equal to y

OpenStudy (anonymous):

okay we can write this eq like this 9x^2-3x-3x+1=0 x=1/3 only one answer as it has real and equal roots for x intercept put y=0 it will give the same i believe :P

OpenStudy (anonymous):

no i have that backwards, its x=-1/3

OpenStudy (anonymous):

it says types answers using radicals as needed

OpenStudy (anonymous):

he's right except its a negative -1/3

OpenStudy (anonymous):

no its not a negative its positive

OpenStudy (anonymous):

9x^2-3x breaks down to 3x(3x-1) so take (3x-1)(3x-1)

OpenStudy (anonymous):

so its positive 3x-1=0 x=1/3

OpenStudy (anonymous):

x=-1/3

OpenStudy (anonymous):

3x-1=0 3x=1 x=1/3

OpenStudy (anonymous):

express in terms of i

OpenStudy (anonymous):

yea, but dont you have top take the opposite of that to equal 0

OpenStudy (anonymous):

yea he's right

OpenStudy (anonymous):

no radicals needed

OpenStudy (anonymous):

please someone give me a straight answer this is my final exam limited time here

OpenStudy (anonymous):

x=1/3

OpenStudy (anonymous):

ok now find x-intercepts

OpenStudy (anonymous):

there is only one x intercept and it's 1/3

OpenStudy (anonymous):

the answer says to type as an ordered pair

OpenStudy (anonymous):

I don't know, there is nothing indicating that there's a y coordinate

OpenStudy (anonymous):

\[-(-6)\pm \sqrt{(-6)^2-4(9)(1)}\] ----------------------------------------- 2(9)

OpenStudy (anonymous):

quadradic formula lol

OpenStudy (anonymous):

6+\[6\pm \sqrt{0}\] ------------------ 18

OpenStudy (anonymous):

Simplifying 9x2 + -6x + 1 = 0 Reorder the terms: 1 + -6x + 9x2 = 0 Solving 1 + -6x + 9x2 = 0 Solving for variable 'x'. Factor a trinomial. (1 + -3x)(1 + -3x) = 0 Subproblem 1Set the factor '(1 + -3x)' equal to zero and attempt to solve: Simplifying 1 + -3x = 0 Solving 1 + -3x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -3x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -3x = 0 + -1 -3x = 0 + -1 Combine like terms: 0 + -1 = -1 -3x = -1 Divide each side by '-3'. x = 0.3333333333 Simplifying x = 0.3333333333 Subproblem 2Set the factor '(1 + -3x)' equal to zero and attempt to solve: Simplifying 1 + -3x = 0 Solving 1 + -3x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -3x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -3x = 0 + -1 -3x = 0 + -1 Combine like terms: 0 + -1 = -1 -3x = -1 Divide each side by '-3'. x = 0.3333333333 Simplifying x = 0.3333333333Solutionx = {0.3333333333, 0.3333333333}

OpenStudy (anonymous):

therefore x = 1/3

OpenStudy (anonymous):

im sry but that is not the correct format

OpenStudy (anonymous):

you are correct, x=1/3 (1/3,0)

OpenStudy (anonymous):

:)